OpenStudy (anonymous):

does the integral x/(5-x)^(1/2) dx that goes from 1 to 5 converge and if so what is the value and how do you find it?

OpenStudy (anonymous):

\[\Large \int_1^5 \frac{x}{\sqrt{5-x}}dx?\]

OpenStudy (anonymous):

yes that's what it looks like...so does it converge? and how do you show that it converges?

OpenStudy (anonymous):

well you do this: \[\Large \lim_{t \rightarrow 5} \int_1^t \frac{x}{\sqrt{5-x}}dx\]

OpenStudy (anonymous):

yup i have that so far :)

OpenStudy (anonymous):

i am thinking trigonometric substitution for this...

OpenStudy (anonymous):